Two long rectilinear conductors are arranged parallel to each other at a

distance of 10 cm. Currents of I1 =I2 =5 A flow through the conductors in

opposite directions. Find the magnitude and direction of the magnetic

field intensity at a point 10 cm from each conductor.


1
Expert's answer
2021-06-18T11:23:32-0400

"B=\frac{\mu_0}{2\pi}\frac IR,"

"B=\frac{2\cdot 5\cdot 10^{-7}}{0.1}=10^{-5}~\text T."